# In-Context Learning as Nonparametric Conditional Probability Estimation: Risk Bounds and Optimality

### v1

@@ 4 The Risk Bounds for In-Context Learning
Let $\alpha\geq 0$ and $\mathcal{H}$ be a function class. A function $\mathbf{p}=(p_{1},\cdots,p_{K})\in\mathcal{H}$ is said to be $\alpha$-small value bounded (or $\alpha$-SVB) if there exists some constant $C>0$ such that for any $k\in[K]$ and $t\in(0,1]$, the following holds:
@@
$\mathcal{G}(\beta,Q):=\left\{\mathbf{p}_{0}=(p_{1}^{0},\dots,p_{K}^{0}):[0,1]^{d}\to\mathcal{S}^{K},\;p_{k}^{0}\in C^{\beta}([0,1]^{d},Q)\text{ for all }k\in[K]\right\},$ (7)
@@
$\log(\mathcal{F}):=\left\{\log\mathbf{p}=(\log p_{1},\dots,\log p_{K}):\mathbf{p}\in\mathcal{F}\right\}.$
Definition 4 ($\delta$-Covering Number and $\delta$-Entropy)
Let $(\mathcal{V},d)$ be a pseudometric space with metric $d(\cdot,\cdot)$ and $\mathcal{F}$ be a subset of $\mathcal{V}$. A finite set $\mathcal{F}(\delta)$ is called a $\delta$-cover (or $\delta$-net) of $\mathcal{F}$ if, for all $\mathbf{f}\in\mathcal{F}$, there exists $\mathbf{g}\in\mathcal{F}(\delta)$ such that $d(\mathbf{f},\mathbf{g})\leq\delta$. Moreover, the $\delta$-covering number of $\mathcal{F}$ with respect to the metric $d$ is the smallest cardinality of such a $\delta$-cover $\mathcal{F}(\delta)$, denoted by $\mathcal{N}(\delta,\mathcal{F},d)$. The quantity $\log\mathcal{N}(\delta,\mathcal{F},d)$ is referred to as the $\delta$-entropy of $\mathcal{F}$ with respect to the metric $d$.
Definition 5 (Empirical and Uniform Supremum Covering Number)
Given $\mathcal{Z}_{T}=\{\mathbf{Z}^{(1)},\ldots,\mathbf{Z}^{(T)}\}\subset\mathbb{R}^{d}$, we define the pseudometric $d=\|\cdot\|_{\infty,\mathcal{Z}_{T}}$ as
$d(\mathbf{f},\mathbf{g}):=\max_{1\leq t\leq T}\max_{1\leq k\leq K}\left|\mathbf{f}_{k}(\mathbf{Z}^{(t)})-\mathbf{g}_{k}(\mathbf{Z}^{(t)})\right|.$
The corresponding metric covering number $\mathcal{N}(\delta,\mathcal{F},\|\cdot\|_{\infty,\mathcal{Z}_{T}})$ is referred as the empirical sup-norm covering number. For a fixed $T$, the largest sup-norm covering number over all $\mathcal{Z}_{T}\subset\mathbb{R}^{d}$ is referred to as the uniform empirical sup-norm covering number:
@@
Let $\mathcal{F}$ be a class of conditional class probability functions $\mathbf{p}=(p_{1},\dots,p_{K})$ used to estimate conditional class probabilities. Suppose that for any $\mathbf{p}\in\mathcal{F}$, there exist two constants $C_{1}>0$ and $0<C_{2}<1$ such that $p_{k}(\mathbf{Z})\geq C_{1}/T^{C_{2}}$ for all $k\in[K]$ and $\mathbf{Z}\in[0,1]^{d}$.
Assumption 2 requires all estimated probabilities to be bounded away from zero. This avoids the singularity of the KL divergence and enables the use of empirical covering numbers to control the complexity of $\log\mathcal{F}$ in the oracle inequality.
@@
Let $\mathcal{F}$ be any class of conditional class probability functions that satisfies Assumption 2, and let $\widehat{\mathbf{p}}\in\mathcal{F}$ be any estimator of $\mathbf{p}_{0}\in\mathcal{S}^{K}$. For any truncation level $B\geq 2$ and task number $T$, the expected excess risk satisfies
$R_{B}(\mathbf{p}_{0},\widehat{\mathbf{p}})\leq C\left(\inf_{\mathbf{p}\in\mathcal{F}}R(\mathbf{p}_{0},\mathbf{p})+\Delta_{T}(\mathbf{p}_{0},\widehat{\mathbf{p}})+\frac{B^{2}\log(\mathcal{N}_{T})\,\log^{2}T}{T}\right),$ (8)
@@
Theorem 7 provides a general oracle inequality for any function class $\mathcal{F}$ satisfying Assumption 2. This result is independent of the specific architecture used to model the conditional class probabilities. In the subsequent analysis, we instantiate $\mathcal{F}$ as a class of neural networks that estimate conditional class probabilities from a given prompt, as defined in (3.4).
@@
To derive the ICL risk, we focus on two key components: the approximation error $\inf_{\mathbf{p}\in\mathcal{F}}R(\mathbf{p}_{0},\mathbf{p})$ and the uniform empirical covering entropy $\mathcal{N}_{T}$ in the oracle inequality (8). We first establish the approximation capacity of transformer networks for functions in the Hölder class $\mathcal{G}(\beta,Q)$ in the following lemma.
@@
Let $0<\varepsilon<1$, $\beta>0$, and $Q>0$. For every function $\mathbf{p}_{0}\in\mathcal{G}(\beta,Q)$, there exists a transformer network $\boldsymbol{\phi}\in\mathcal{T}(M,H,D,W_{1},B_{1},S_{1},\gamma)$ with
@@
Lemma 8 indicates that transformer networks can uniformly approximate any function in the Hölder class $\mathcal{G}(\beta,Q)$ with explicit control over architectural complexity. Similar approximation results have been established in related settings, such as Theorem 8 in Jiao et al. (2024). Its proof builds on the observation by Gurevych et al. (2022) that transformer architectures, such as deep ReLU networks, can represent multiplication operations ans thus approximate polynomials. Since polynomials can can approximate continuous functions arbitrarily well, transformer networks can effectively approximate Hölder-smooth targets. Without loss of generality, we assume the output of the approximating transformer lies in $[0,1]$; otherwise, a projection layer can be applied without increasing the error, as the target functions $p_{k}^{0}$ are already supported on this interval.
Next, we apply Lemma 8 to approximate the conditional class probabilities in multiclass classification. Specifically, we consider the neural network class $\mathcal{F}(M,H,D,W_{1},B_{1},S_{1},\gamma,L,\\ W,S)$ defined in (3.4), which is constructed to approximate the ground-truth conditional class probabilities $\mathbf{p}_{0}\in\mathcal{G}(\beta,Q)$ in the $\|\cdot\|_{\infty}$-norm, while ensuring that the output probability vectors are uniformly bounded away from zero. This is achieved by composing a transformer encoder with an MLP followed by a softmax transformation. The MLP processes the representations produced by the transformer encoder and computes a vector of logits, which are then mapped by the softmax function to class probability vectors. The resulting outputs are uniformly bounded away from zero and thus satisfy Assumption 2. A similar construction is used in Bos and Schmidt-Hieber (2022).
This transformer-MLP hybrid architecture is also motivated by its practical effectiveness. Transformers provide expressive, context-aware representations through self-attention mechanisms, while the MLP acts as an adaptive task-specific output head. Such modular designs are widely used in contemporary deep learning models and have demonstrated strong generalization performance across diverse applications. For simplicity, we henceforth denote the neural network classes $\mathcal{T}(M,H,D,W_{1},B_{1},S_{1},\gamma)$ defined in (3.3), $\mathcal{F}_{\mathrm{id}}(L,W,S)$ defined in (5), and $\mathcal{F}(M,H,D,W_{1},B_{1},S_{1},\gamma,L,W,S)$ defined in (3.4) as $\mathcal{T}$, $\mathcal{F}_{\mathrm{id}}$, and $\mathcal{F}$, respectively.
@@
Fix $\beta>0$ and $Q>0$. For any $\mathbf{p}_{0}\in\mathcal{G}(\beta,Q)$ and any $\varepsilon>0$, there exists a neural network $\mathbf{\tilde{q}}=(\tilde{q}_{1},\cdots,\tilde{q}_{K})\in\mathcal{F}$ with
@@
$\|{\tilde{q}}_{k}-{p}^{0}_{k}\|_{\infty}<2\varepsilon(K+1),\quad{\tilde{q}}_{k}(\mathbf{Z})\geq\frac{\varepsilon}{2(K+1)}$
hold for all $k\in\{1,2,\dots,K\}$ and $\mathbf{Z}\in[0,1]^{d}$.
Theorem 9 shows that the conditional probability $\mathbf{p}_{0}$ can be uniformly approximated, with arbitrarily small error, by a neural network from $\mathcal{F}$. Its proof is provided in Appendix B. Based on this result, we derive a non-asymptotic upper bound for the minimal risk $\inf_{\mathbf{p}\in\mathcal{F}}R(\mathbf{p}_{0},\mathbf{p})$, under uniform approximation and margin conditions.
Remark 10
The neural network $\mathbf{\tilde{q}}$ constructed in Theorem 9 inherently satisfies a natural lower bound on its output for every function $\mathbf{p}_{0}\in\mathcal{G}(\beta,Q)$. Specifically, for any $\varepsilon>0$, the network’s outputs are bounded below by $\frac{\varepsilon}{2(K+1)}$, which is an intrinsic property of the architecture. This lower bound aligns with Assumption 2. In particular, by selecting an appropriate $\varepsilon$, the neural network $\mathbf{\tilde{q}}$ can be guaranteed to meet the lower bound condition on the probability functions, as specified by Assumption 2.
@@
Assume $\mathbf{p}_{0}\in\mathcal{G}_{\alpha}(\beta,Q)$ for some $\alpha\in[0,1]$. Let $\mathbf{p}=(p_{1},\ldots,p_{K}):[0,1]^{d}\to\mathcal{S}^{K}$ be an approximation of $\mathbf{p}_{0}$ such that
$\|\mathbf{p}-\mathbf{p}_{0}\|_{\infty}\leq 2(K+1)\varepsilon,\quad\min_{1\leq k\leq K}\inf_{\mathbf{Z}\in[0,1]^{d}}p_{k}(\mathbf{Z})\geq\frac{\varepsilon}{2(K+1)}.$
@@
If we choose $\varepsilon=C^{-1}T^{-\frac{\beta}{(1+\alpha)\beta+d}}K^{-\frac{(3+\alpha)\beta}{(1+\alpha)\beta+d}}$ with $C>0$ being a sufficiently small constant, then the risk $\inf_{\mathbf{p}\in\mathcal{F}}R(\mathbf{p}_{0},\mathbf{p})$ is bounded as
Theorem 11 establishes a near-optimal convergence rate for in-context classification, recovering the classical nonparametric minimax bound when $\alpha=1$. The proof leverage $f$-divergence theory, bounding the KL divergence via the $\chi^{2}$-divergence (Nguyen et al., 2009; Cai and Lim, 2022). More details can be found in Appendix C.
Next, we deal with the uniform empirical covering entropy $\log(\mathcal{N}_{T})$ in the Oracle Inequality. Recall that
$\log(\mathcal{N}_{T})=\mathcal{V}_{\infty}\Big{(}\frac{1}{T},\log\mathcal{F},T\Big{)}=\sup_{\mathcal{Z}_{T}}\log\mathcal{N}\Big{(}\frac{1}{T},\log\mathcal{F},\|\cdot\|_{\infty,\mathcal{Z}_{T}}\Big{)}.$
The following lemma provides an upper bound on the uniform empirical covering entropy $\log(\mathcal{N}_{T})$ of the class $\log(\mathcal{F})$ under the empirical sup-norm.
@@
For any $\delta>0$, and for sufficiently small $\delta$, the following inequality holds:
$\mathcal{V}_{\infty}\left(\delta,\log\mathcal{F},T\right)\leq\mathcal{V}_{\infty}\left(\tfrac{\delta}{4KL_{1}},\,\mathcal{T},\,T\right)+\mathcal{V}\left(\tfrac{\delta}{4K},\,\mathcal{F}_{\mathrm{id}},\,\|\cdot\|_{\infty}\right),$
@@
The proof of Lemma 12 is provided in Appendix D. The argument decomposes $\log\mathcal{F}$ into a composition of the transformer class $\mathcal{T}$ and the MLP class $\mathcal{F}_{\mathrm{id}}$, and then applies standard covering number bounds for Lipschitz compositions to control the resulting entropy.
The next result provides an upper bound on the uniform empirical covering entropy of the transformer class under the empirical sup-norm $\|\cdot\|_{\infty,\mathcal{Z}_{T}}$, where $\mathcal{Z}_{T}=\{\mathbf{Z}^{(1)},\dots,\mathbf{Z}^{(T)}\}$.
@@
For any $\delta_{1}>0$, the uniform empirical covering entropy of the transformer class $\mathcal{T}$ under the empirical sup-norm satisfies
@@
where $\mathcal{T}=\mathcal{T}(M,H,D,W_{1},B_{1},S_{1},\gamma)$.
@@
For any $\delta_{2}>0$, the covering entropy of the function class $\mathcal{F}_{\mathrm{id}}$ under the sup-norm $\|\cdot\|_{\infty}$ satisfies the following upper bound
@@
Lemma 14 gives an upper bound on the uniform covering entropy of the class $\mathcal{F}_{\mathrm{id}}(L,W,S)$, which consists of $S$-sparse ReLU networks with depth $L$, width $W$, and output dimension $K$, measured under the sup-norm over the input domain. Its proof is provided in Appendix F.
@@
Assume $\mathbf{p}_{0}\in\mathcal{G}_{\alpha}(\beta,Q)$ for some $\alpha\in[0,1]$. Let $\hat{\mathbf{p}}$ be an estimator of $\mathbf{p}_{0}$, taking values in the function class $\mathcal{F}(M,H,D,W_{1},B_{1},S_{1},\gamma,L,W,S)$, where the parameters of the class scale with the precision parameter $\varepsilon>0$ as follows:
@@
$R_{B}\left(\mathbf{p}_{0},\hat{\mathbf{p}}\right)\lesssim B^{2}T^{-\frac{(1+\alpha)\beta}{(1+\alpha)\beta+d}}K^{4+\alpha}\log^{5}\left(KT\right).$
@@
$R_{B}\left(\mathbf{p}_{0},\hat{\mathbf{p}}\right)\lesssim B^{2}T^{-\frac{2\beta}{2\beta+d}}K^{5}\log^{5}\left(KT\right).$
Theorem 15 establishes that the ICL risk attains the minimax-optimal convergence rate (up to logarithmic factors) when $\alpha=1$. This result highlights the importance of pretraining task diversity, as observed by Raventós et al. (2023). Our setup corresponds to the task-scaling regime formalized by Abedsoltan et al. (2024), where the risk decreases as the number of pretraining tasks $T$ increases, while the context length remains fixed. In our multiclass setting, the achievable risk is fundamentally determined by the number of pretraining tasks $T$. Since large language models are trained on virtually unlimited tasks, our bound provides a theoretical explanation for their strong in-context learning performance—even with only a small number of in-context examples at inference time.
Proof of Theorem 15. The proof is based on the oracle inequality (8), which decomposes the error into an approximation error $\inf_{\mathbf{p}\in\mathcal{F}}R(\mathbf{p}_{0},\mathbf{p})$, controlled by Theorem 11, and a complexity term $(\log(\mathcal{N}_{T})\log^{2}T)/T$, bounded using Lemmas 12, 13, and 14, with $\delta=1/T$. The approximation scale $\varepsilon=C^{-1}T^{-\frac{\beta}{(1+\alpha)\beta+d}}K^{-\frac{(3+\alpha)\beta}{(1+\alpha)\beta+d}}$ is chosen to balance the approximation error and the entropy term, reflecting the trade-off between network expressiveness and statistical complexity. These bounds together yields our result. $\blacksquare$

### v2

@@ 4 The Risk Bounds for In-Context Learning
Let $\alpha\geq 0$ and $\mathcal{H}$ be a function class. A function $\mathbf{p}=(p_{1},\dots,p_{K})\in\mathcal{H}$ is said to be $\alpha$-small value bounded (or $\alpha$-SVB) if there exists some constant $C>0$ such that for any $k\in[K]$ and $t\in(0,1]$, the following holds:
@@
$\mathcal{G}(\beta,Q):=\left\{\mathbf{p}_{0}=(p_{1}^{0},\dots,p_{K}^{0})^{\top}:[0,1]^{d}\to\mathcal{S}^{K},\;p_{k}^{0}\in C^{\beta}([0,1]^{d},Q)\text{ for all }k\in[K]\right\},$ (7)
@@
$\log(\mathcal{F}):=\left\{\log\mathbf{p}=(\log p_{1},\dots,\log p_{K})^{\top}:\mathbf{p}\in\mathcal{F}\right\}.$
Definition 4 ($\delta$-Covering Number and Metric Entropy)
Let $(\mathcal{V},d)$ be a pseudometric space with metric $d(\cdot,\cdot)$ and $\mathcal{F}$ be a subset of $\mathcal{V}$. A finite set $\mathcal{F}(\delta)$ is called a $\delta$-cover (or $\delta$-net) of $\mathcal{F}$ if, for all $\mathbf{f}\in\mathcal{F}$, there exists $\mathbf{g}\in\mathcal{F}(\delta)$ such that $d(\mathbf{f},\mathbf{g})\leq\delta$. Moreover, the $\delta$-covering number of $\mathcal{F}$ with respect to the metric $d$ is the smallest cardinality of such a $\delta$-cover $\mathcal{F}(\delta)$, denoted by $\mathcal{N}(\delta,\mathcal{F},d)$. The quantity $\log\mathcal{N}(\delta,\mathcal{F},d)$ is referred to as the metric entropy of $\mathcal{F}$ at scale $\delta$ with respect to the metric $d$.
Definition 5 (Uniform Empirical Covering Number and Entropy)
For any given $\mathcal{Z}_{T}=\{\mathbf{Z}^{(1)},\ldots,\mathbf{Z}^{(T)}\}\subset\mathbb{R}^{d}$, we define the pseudometric $d(\mathbf{f},\mathbf{g})=\|\mathbf{f}-\mathbf{g}\|_{\infty,\mathcal{Z}_{T}}$ as
$d(\mathbf{f},\mathbf{g})=\|\mathbf{f}-\mathbf{g}\|_{\infty,\mathcal{Z}_{T}}:=\max_{1\leq t\leq T}\max_{1\leq k\leq K}\left|\mathbf{f}_{k}(\mathbf{Z}^{(t)})-\mathbf{g}_{k}(\mathbf{Z}^{(t)})\right|.$
The corresponding metric covering number $\mathcal{N}(\delta,\mathcal{F},\|\cdot\|_{\infty,\mathcal{Z}_{T}})$ is referred as the empirical covering number with respect to $\|\cdot\|_{\infty,\mathcal{Z}_{T}}$. For a fixed $T$, the largest covering number over all $\mathcal{Z}_{T}\subset\mathbb{R}^{d}$ is referred to as the uniform empirical covering number:
@@
The uniform empirical entropy is defined as $\mathcal{V}_{\infty}(\delta,\mathcal{F},T):=\log\mathcal{N}_{\infty}(\delta,\mathcal{F},T)$.
@@
Let $\mathcal{F}$ be a class of conditional class probability functions $\mathbf{p}=(p_{1},\dots,p_{K})^{\top}$ used to estimate conditional class probabilities. Suppose that for any $\mathbf{p}\in\mathcal{F}$, there exist two constants $C_{1}>0$ and $0<C_{2}<1$ such that $p_{k}(\mathbf{Z})\geq C_{1}/T^{C_{2}}$ for all $k\in[K]$ and $\mathbf{Z}\in[0,1]^{d}$.
Assumption 2 requires all estimated probabilities to be bounded away from zero. This avoids the singularity of the KL divergence and enables the use of empirical covering numbers to control the complexity of $\log(\mathcal{F})$ in the oracle inequality. For the neural network class $\mathcal{F}(M,H,D,W_{1},B_{1},S_{1},\gamma,L,W,S)$ defined in (3.4), this can be achieved by adding additional layers to the MLP architecture of the network. More details on this issue are provided in Remarks 11 and 30.
@@
Let $\mathcal{F}$ be any class of conditional class probability functions that satisfies Assumption 2, and let $\widehat{\mathbf{p}}\in\mathcal{F}$ be an estimator of $\mathbf{p}_{0}\in\mathcal{S}^{K}$. For any truncation level $B\geq 2$ and task number $T$, the expected excess risk satisfies
$R_{B}(\mathbf{p}_{0},\widehat{\mathbf{p}})\lesssim\inf_{\mathbf{p}\in\mathcal{F}}R(\mathbf{p}_{0},\mathbf{p})+\Delta_{T}(\mathbf{p}_{0},\widehat{\mathbf{p}})+\frac{B^{2}\log(\mathcal{N}_{T})\,\log^{2}T}{T},$ (8)
@@
The proof of Theorem 7 is provided in Appendix A. It provides a general oracle inequality for any function class $\mathcal{F}$ satisfying Assumption 2. This result is independent of the specific architecture used to model the conditional class probabilities.
@@
Let $0<\varepsilon<1$, $\beta>0$, and $Q>0$. For every function $\mathbf{p}_{0}\in\mathcal{G}(\beta,Q)$, there exists a transformer network $\boldsymbol{\phi}\in\mathcal{T}$ with
@@
Lemma 8 indicates that transformer networks can uniformly approximate any function in the Hölder class $\mathcal{G}(\beta,Q)$ with explicit control over architectural complexity. Similar approximation results have been established in related settings, such as Theorem 8 in Jiao et al. (2024). Its proof builds on the observation by Gurevych et al. (2022) that transformer architectures, similar to deep ReLU networks, are capable of representing multiplication operations and therefore can approximate polynomials. Since polynomials can approximate continuous functions with an arbitrarily accuracy, transformer networks can effectively approximate Hölder-smooth targets. Without loss of generality, we assume the output of the approximating transformer lies in $[0,1]$; otherwise, a projection layer can be applied without increasing the error, as the target functions $p_{k}^{0}$ are already supported on this interval.
Next, we apply Lemma 8 to approximate the conditional class probabilities in multiclass classification. Specifically, we consider the neural network class $\mathcal{F}$ defined in (3.4) to approximate the true conditional class probabilities $\mathbf{p}_{0}\in\mathcal{G}(\beta,Q)$ in the $\|\cdot\|_{\infty}$-norm, while ensuring that the output probabilities are uniformly bounded away from zero. This is achieved by composing a transformer encoder with a MLP, followed by a softmax transformation. The detailed construction of the approximating network in $\mathcal{F}$ is provided in the proof of Theorem 9 in Appendix B.
This transformer-MLP hybrid architecture is also motivated by its practical effectiveness. Transformers provide expressive, context-aware representations through self-attention mechanisms, while the MLP acts as an adaptive task-specific output head. Such modular designs are widely used in contemporary deep learning models and have demonstrated strong generalization performance across diverse applications.
@@
Fix $\beta>0$ and $Q>0$. For any $\mathbf{p}_{0}\in\mathcal{G}(\beta,Q)$ and $\varepsilon\in(0,1)$, there exists a neural network $\mathbf{\tilde{q}}=(\tilde{q}_{1},\cdots,\tilde{q}_{K})\in\mathcal{F}$ with
@@
$\|{\tilde{q}}_{k}-{p}^{0}_{k}\|_{\infty}\leq 2\varepsilon(K+1),\quad{\tilde{q}}_{k}(\mathbf{Z})\geq\frac{\varepsilon}{2K}$
hold for all $k\in[K]$ and $\mathbf{Z}\in[0,1]^{d}$.
Theorem 9 shows that the conditional probability $\mathbf{p}_{0}\in\mathcal{G}(\beta,Q)$ can be uniformly approximated, with arbitrarily small error, by a neural network of $\mathcal{F}$. Its proof is provided in Appendix B. Based on this result, we derive a non-asymptotic upper bound for the approximation error $\inf_{\mathbf{p}\in\mathcal{F}}R(\mathbf{p}_{0},\mathbf{p})$.
@@
Assume $\mathbf{p}_{0}\in\mathcal{G}_{\alpha}(\beta,Q)$ for some $\alpha\in[0,1]$. Let $\mathbf{p}=(p_{1},\ldots,p_{K})^{\top}:[0,1]^{d}\to\mathcal{S}^{K}$ be an approximation of $\mathbf{p}_{0}$ such that
$\|\mathbf{p}-\mathbf{p}_{0}\|_{\infty}\leq 2(K+1)\varepsilon,\quad\min_{1\leq k\leq K}\inf_{\mathbf{Z}\in[0,1]^{d}}p_{k}(\mathbf{Z})\geq\frac{\varepsilon}{2K}.$
@@
If we choose $\varepsilon=C^{-1}T^{-\frac{\beta}{(1+\alpha)\beta+d}}K^{-\frac{(3+\alpha)\beta}{(1+\alpha)\beta+d}}$ for some constant $C>0$, then
$\inf_{\mathbf{p}\in\mathcal{F}}R(\mathbf{p}_{0},\mathbf{p})\lesssim K^{4+\alpha}T^{-\frac{(1+\alpha)\beta}{(1+\alpha)\beta+d}}\log(KT).$
Theorem 10 establishes a near-optimal convergence rate for in-context classification, recovering the classical nonparametric minimax bound when $\alpha=1$. The proof leverage $f$-divergence theory, bounding the KL divergence via the $\chi^{2}$-divergence (Nguyen et al., 2009; Cai and Lim, 2022). More details about the proof can be found in Appendix C.
Remark 11
According to the proof of Theorem 9 and Remark 30 in Appendix B, we can add additional ReLU layers to the MLP architecture of the network in $\mathcal{F}$ to ensure that $p_{k}(\mathbf{Z})\geq\varepsilon/(2K)$ for any network $\mathbf{p}=(p_{1},\dots,p_{K})^{\top}\in\mathcal{F}$ and $\varepsilon\in(0,1)$. Consequently, without loss of generality, we require that every network in $\mathcal{F}$ has a uniform, strictly positive lower bound $\varepsilon/(2K)$ across all its coordinates. Furthermore, if we choose $\varepsilon=C^{-1}T^{-\frac{\beta}{(1+\alpha)\beta+d}}K^{-\frac{(3+\alpha)\beta}{(1+\alpha)\beta+d}}$, it is readily seen that the function class $\mathcal{F}$ satisfies Assumption 2. This allows us to apply the oracle inequality (Theorem 7) to derive the expected excess risk for transformer-based classifiers.
Next, we deal with the uniform empirical entropy $\log(\mathcal{N}_{T})$ in the oracle inequality. Recall that
$\log(\mathcal{N}_{T})=\mathcal{V}_{\infty}\Big{(}\frac{1}{T},\log(\mathcal{F}),T\Big{)}=\sup_{\mathcal{Z}_{T}}\log\mathcal{N}\Big{(}\frac{1}{T},\log(\mathcal{F}),\|\cdot\|_{\infty,\mathcal{Z}_{T}}\Big{)},$
where $\mathcal{Z}_{T}=\{\mathbf{Z}^{(1)},\ldots,\mathbf{Z}^{(T)}\}\subset\mathbb{R}^{d}$. The following lemma provides an upper bound on the uniform empirical entropy $\log(\mathcal{N}_{T})$ of the function class $\log(\mathcal{F})$.
@@
For any $\delta>0$, the following inequality holds:
$\mathcal{V}_{\infty}\left(\delta,\log(\mathcal{F}),T\right)\leq\mathcal{V}_{\infty}\left(\tfrac{\delta}{4KL_{1}},\,\mathcal{T},\,T\right)+\mathcal{V}\left(\tfrac{\delta}{4K},\,\mathcal{F}_{\mathrm{id}},\,\|\cdot\|_{\infty}\right),$
@@
The proof of Lemma 12 is provided in Appendix D. The argument decomposes $\log(\mathcal{F})$ into a composition of the transformer class $\mathcal{T}$ and the MLP class $\mathcal{F}_{\mathrm{id}}$, and then applies standard covering number bounds for Lipschitz compositions to control the resulting entropy. The next result provides an upper bound on the uniform empirical entropy of the transformer class.
@@
For any $\delta_{1}>0$, the uniform empirical entropy of the transformer class $\mathcal{T}$ satisfies
@@
where $\mathcal{T}=\mathcal{T}(M,H,D,W_{1},B_{1},S_{1},\gamma)$ is given in (3.3).
@@
For any $\delta_{2}>0$, the metric entropy of the function class $\mathcal{F}_{\mathrm{id}}$ satisfies the following upper bound
@@
[…]
@@
Assume $\mathbf{p}_{0}\in\mathcal{G}_{\alpha}(\beta,Q)$ for some $\alpha\in[0,1]$. Let $\hat{\mathbf{p}}$ be an estimator of $\mathbf{p}_{0}$, taking values in the function class $\mathcal{F}(M,H,D,W_{1},B_{1},S_{1},\gamma,L,W,S)$, where the parameters of this class scale with the precision parameter $\varepsilon>0$ as follows:
@@
$R_{B}\left(\mathbf{p}_{0},\hat{\mathbf{p}}\right)\lesssim B^{2}K^{4+\alpha}T^{-\frac{(1+\alpha)\beta}{(1+\alpha)\beta+d}}\log^{7}\left(KT\right).$
@@
$R_{B}\left(\mathbf{p}_{0},\hat{\mathbf{p}}\right)\lesssim B^{2}K^{5}T^{-\frac{2\beta}{2\beta+d}}\log^{7}\left(KT\right).$
Theorem 15 establishes that the upper bound of the ICL risk attains the minimax optimal rate $T^{-2\beta/(2\beta+d)}$ (up to logarithmic factors) when $\alpha=1$. This result highlights the importance of pretraining task diversity, as observed by Raventós et al. (2023). Our setup corresponds to the task-scaling regime formalized by Abedsoltan et al. (2024), where the risk decreases as the number of pretraining tasks $T$ increases, while the context length $N$ remains fixed. In our multiclass setting, the achievable risk is fundamentally determined by the number of pretraining tasks $T$. Since large language models are trained on virtually unlimited tasks, our bound provides a theoretical explanation for their strong in-context learning performance, even with only a small number of in-context examples at inference time.
[…]